OpenStudy (anonymous):

....which expression is equal to the expression below ? ( 6 *9 )^5 a. 6^5*9^5 b.6^5 + 9^5 c. 15 ^5 d. 5*(6*9) ? help ???

OpenStudy (wasiqss):

d is rite

OpenStudy (anonymous):

Wasiq Moderators said your supposed to explain not just give answers thats not what OS is for.

jimthompson5910 (jim_thompson5910):

Use the idea that (x*y)^z = x^z * y^z to get (6*9)^5 = 6^5 * 9^5

Directrix (directrix):

a. 6^5*9^5 ----------- ( 6 *9 )^5